# AAA-Gaussians: Anti-Aliased and Artifact-Free 3D Gaussian Rendering

### v1

@@ 4.2 View-Consistent Rendering
$\frac{1}{2}{\times}$ MCMC 28.98 0.941 0.099 21.15 0.774 0.171
@@
Popping artifacts due to culling issues are difficult to evaluate quantitatively, as ground truth data cannot be generated from existing evaluation views. An example for this kind of popping is shown in Fig. 1. For more examples, please see the supplemental video.
@@ 4.3 Performance Timings
We evaluate the runtime performance of our components and MCMC in Tab. 5, using an NVIDIA RTX 4090 with timings averaged over an interpolated camera path across all available camera poses. Our method is only slightly slower than MCMC (with standard 3DGS bounding) and even outperforms it on the Tanks and Temples dataset. When removing culling, performance drops significantly, as the hierarchical sort heavily relies on culling to reduce its sorting overhead. Disabling hierarchical sorting improves speed beyond MCMC but introduces noticeable popping artifacts. Lastly, disabling our 3D evaluation results in similar or worse performance, indicating that our 3D evaluation is as fast or even faster than 2D splat variants.
@@ 5 Conclusion, Limitations, And Future Work
In this work, we addressed the limitations of current 3D Gaussian Splatting methods and made several key contributions to enable fast, artifact-free rendering of 3D Gaussians. Our method introduces a novel 3D smoothing filter that effectively removes aliasing artifacts when evaluating Gaussians in 3D, along with a stable 3D bounding and culling approach that performs consistently across various viewing scenarios. We thoroughly evaluated the effectiveness of our components, showing that our method is robust to out-of-distribution camera views while maintaining standard image metrics on par with the current 3DGS state-of-the-art. To our knowledge, we are the only rasterization-based approach capable of delivering artifact-free rendering of 3D Gaussians, with framerates exceeding 100 FPS on consumer-grade hardware.
@@ Appendix A Derivation of Amplitude Scaling Factor
  • $\sigma_{11}\in\mathbb{R}$ is the $(1,1)$ entry,
@@
Using the standard formula for the inverse of a block matrix, we can rewrite the reciprocal of the Schur complement as the $(1,1)$ entry of $\mathbf{\Sigma}^{\prime-1}$,
@@ Appendix C Multi-Resolution Evaluation Images
We show an example view of our multi-resolution evaluation in Fig. 7. While Ours is able to retain good image quality at all resolution levels, and correctly dilates and smooths content. In contrast, MCMC and our method without the anti-aliasing 3D smoothing filter exhibit considerable aliasing: content becomes too thick on lower resolution and too thin on higher resolution. This also shows in the inset PSNR values.
@@ Appendix D Per Scene Image Metrics
We provide per-scene image metrics for all our evaluated scenes in Tab. 6.

### v2

@@ 4.2 View-Consistent Rendering
$\frac{1}{2}{\times}$
MCMC 28.98 0.941 0.099 21.15 0.774 0.171
@@
Popping artifacts caused by incorrect culling when moving close to scene content—which our method resolves—are difficult to evaluate quantitatively, as ground truth data cannot be generated from existing evaluation views. An example for this kind of popping is shown in Fig.˜1. For more examples, please see the supplemental video.
@@ 4.3 Performance Timings
We evaluate the runtime performance of our components and MCMC in Tab.˜5, using an NVIDIA RTX 4090 with timings averaged over an interpolated camera path across all available camera poses. Our method is only slightly slower than MCMC (with standard 3DGS bounding) and even outperforms it on the Tanks and Temples dataset. When removing culling, performance drops significantly, as the hierarchical sort heavily relies on culling to reduce its sorting overhead. Disabling hierarchical sorting improves speed beyond MCMC but introduces noticeable popping artifacts. Lastly, disabling our 3D evaluation results in similar or worse performance, indicating that our 3D evaluation is as fast or even faster than 2D splat variants.
@@ 5 Conclusion, Limitations, And Future Work
In this work, we addressed the limitations of current 3D Gaussian Splatting methods and made several key contributions to enable fast, artifact-free rendering of 3D Gaussians. Our method introduces a novel 3D smoothing filter that effectively removes aliasing artifacts when evaluating Gaussians in 3D, along with stable 3D bounding and culling that performs consistently across various viewing scenarios. We thoroughly evaluated the effectiveness of our components, showing that our method is robust to out-of-distribution camera views while maintaining standard image metrics on par with the current 3DGS state-of-the-art. To our knowledge, we are the only rasterization-based approach capable of delivering artifact-free rendering of 3D Gaussians, with framerates exceeding 100 FPS on consumer-grade hardware.
@@
Acknowledgments
This research was supported by the Austrian Science Fund FWF [10.55776/I6663], the German Science Foundation DFG [contract 528364066] and the Alexander von Humboldt Foundation funded by the German Federal Ministry of Research, Technology and Space. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
@@ Appendix A Derivation of Amplitude Scaling Factor
  • $\sigma_{11}\in\mathbb{R}$ is the (1,1)(1,1)( 1 , 1 ) entry,
@@
Using the standard formula for the inverse of a block matrix, we can rewrite the reciprocal of the Schur complement as the (1,1)(1,1)( 1 , 1 ) entry of $\mathbf{\Sigma}^{\prime-1}$,
@@ Appendix C Multi-Resolution Evaluation Images
We show an example view of our multi-resolution evaluation in Fig.˜7. While Ours is able to retain good image quality at all resolution levels, and correctly dilates and smooths content. In contrast, MCMC and our method without the anti-aliasing 3D smoothing filter exhibit considerable aliasing: content becomes too thick on lower resolution and too thin on higher resolution. This also shows in the inset PSNR values.
@@ Appendix D Per Scene Image Metrics
We provide per-scene image metrics for all our evaluated scenes in Tab.˜6.